OpenStudy (anonymous):

The temperature T of a given mass of gas varies inversely with its volume V. The temperature of 90 cm3 of a certain gas is 30°C. What will be the temperature of the gas when it is compressed to a volume of 18 cm3?

OpenStudy (anonymous):

V1T1=V2T2 so possibly 90*30=18*T solve for T

OpenStudy (anonymous):

interesting that they would give you this problem considering a chemistry law called Charles' Law states that they are directly related V/T=V2/T2